# biblio.bib

### v2

@@ Abstract
We describe a class of spin chains with new physical and computational properties. On the physical side, the spin chains give examples of symmetry-protected topological phases that are defined by non-onsite symmetries, i.e. symmetries that are not a tensor product of single-site operators. These phases can be detected by string-order parameters, but notably do not exhibit entanglement spectrum degeneracy. On the computational side, the spin chains represent a new class of states that can be used to deterministically teleport information across long distances, with the novel property that the necessary classical side processing is a non-linear function of the measurement outcomes. We also give examples of states that can serve as universal resources for measurement-based quantum computation, providing the first examples of such resources without entanglement spectrum degeneracy. The key tool in our analysis is a new kind of tensor network representation which we call split-index matrix product states (SIMPS). We develop the basic formalism of SIMPS, compare them to matrix product states, show how they are better equipped to describe certain kinds of non-onsite symmetries including anomalous symmetries, and discuss how they are also well-suited to describing quantum teleportation and constrained spin chains.
@@ 1 Introduction
In this paper, we construct a class of quantum spin chains that have novel physical and computational properties. The key concept that we introduce is a new kind of tensor network representation we call split-index matrix product states (SIMPS) (4). While equivalent to MPS in terms of which states can be efficiently captured, SIMPS offer an alternative description that we argue is essential to understanding the properties of the constructed models. For example, a key physical property of the models is that they possess symmetries which are non-onsite, i.e. not a tensor product of single-site unitaries. Compared to MPS, we show that SIMPS provide a much simpler way to identify and analyze such symmetries. Similarly, SIMPS offer a clearer picture of the computational capabilities of certain models, including ones that have previously appeared in the literature. We also show that SIMPS are well-suited to capturing constrained spin chains, including models exhibiting quantum many-body scars and Hilbert space fragmentation [Moudgalya2022].
@@ 2 The SIMPS representation
As in the case of MPS, this property is stable in that it holds for all $L>L_{1}$ as well, and it allows us to construct an operator that reveals the virtual space of the SIMPS. Define the map,
@@
which acts on $L+2$ spins and is invertible for $L\geq L_{1}$. Then, $B_{L}^{-1}$ serves the purpose of revealing the virtual indices of the SIMPS, which we will use to construct string order parameters later. One could also use this map to construct parent Hamiltonians for SIMPS which may have different properties than they corresponding MPS parent Hamiltonians, but we do not pursue this here.
@@ 2.1 Converting between MPS and SIMPS
First, suppose we have a SIMPS tensor $B^{ij}$. For simplicity, we assume here that $\chi_{i}=\chi$ for all $i$. A similar construction for non-uniform $\chi_{i}$ can be obtained by appropriately padding the matrices $B^{ij}$ with zeros. We can convert the SIMPS to an MPS by cutting $B^{ij}$ in half via singular value decomposition (SVD). Namely, define the matrix $\mathbb{B}=\sum_{i,j,\alpha,\beta}B^{ij}_{\alpha\beta}|i\alpha\rangle\langle j\beta|$ that maps $\mathbb{C}^{d}\otimes\mathbb{C}^{\chi}$ onto itself and use the SVD to write $\mathbb{B}=usv^{\dagger}$ where $u,v$ are isometries mapping from $\mathbb{C}^{d}\otimes\mathbb{C}^{\chi}$ to $\mathbb{C}^{D}$ satisfying $u^{\dagger}u=v^{\dagger}v=\mathbb{1}$ and $s$ is a $D\times D$ diagonal matrix. Therein, $D$ is the rank of $\mathbb{B}$. For compactness, let us absorb $s$ into $v$, defining $w^{\dagger}=sv^{\dagger}$. From this, we define the matrices $A^{i}$ as,
@@
In the generic case, $A^{i}$ is full rank so $P^{i}=\mathbb{1}$ and $B^{ij}=A^{i}$. The cases where the SIMPS representation is distinct from the MPS representation are therefore those for which $A^{i}$ is not full-rank in some physical basis $\{i\}$. This leads us to the following definition of when a MPS “makes a good SIMPS”; i.e. when the SIMPS and MPS representations differ,
@@ 2.2 Fundamental theorem of SIMPS
We now state the SIMPS equivalent of the fundamental theorem of MPS. The theorem describes how two normal SIMPS tensors generating the same state are related. For MPS, the fundamental theorem states that if $A^{i}$ and $B^{i}$ are both normal and generate the same MPS, then there is an invertible matrix $V$ such that $A^{i}=VB^{i}V^{-1}$, i.e. the two tensors are related by a so-called gauge transformation [Perez-Garcia2008]. For injective SIMPS, gauge transformations like $B^{ij}\rightarrow VB^{ij}V^{-1}$ can still be used to obtain new tensors generating the same state. However, it turns out that a larger class of transformations is also possible,
@@ 2.3 Examples of SIMPS
Example 1: The cluster state. First, we consider the 1D cluster state, which is a paradigmatic resource state for MBQC [Raussendorf2001a], and also a simple of example of a state with SPT order [Son2012]. The cluster state can be written as,
@@
It is straightforward to verify that this MPS in normal, so it describes the unique ground state of a gapped Hamiltonian. Both matrices are rank deficient, having rank equal to two. Therefore, using Eq. (9), we can derive the corresponding SIMPS tensor with bond dimension $\chi=2$,
@@ 3 Non-onsite symmetries and SPT phases in SIMPS
In this section, we demonstrate the one context where SIMPS are advantageous over MPS: capturing non-onsite symmetries. First, we review how onsite symmetries are capture in MPS. Then, we show how similar ideas can be used to analyze certain non-onsite symmetries in SIMPS. This includes a general class of anomalous group-like symmetries. We then define a class of SIMPS possessing non-trivial SPT order protected by non-onsite symmetries and characterize their physical properties.
@@ 3.1 SPT order in MPS
If there is a set of symmetries $u_{g}$ that form a representation of a group $G$ such that $u_{g}u_{h}=u_{gh}$, then the corresponding operators $V_{g}$ that appear in Eq. (21) form a projective representation, $V_{g}V_{h}=\omega(g,h)V_{gh}$ for some phase factors $\omega(g,h)$ [Schuch2011]. If $G$ is finite abelian, which we always assume from now on, we can instead characterize the representation by the relation $V_{g}V_{h}=\Omega(g,h)V_{h}V_{g}$ where $\Omega(g,h)=\omega(g,h)\omega(h,g)^{-1}$. The phases $\Omega(g,h)$ uniquely label the SPT phase to which the MPS belongs [Berkovich1998]. If these phases are not all equal to one, then the MPS belongs to a non-trivial SPT phase protected by $G$ [Pollmann2010, Chen2011, Schuch2011].
@@ 3.3 Anomalous symmetries
Non-onsite symmetries are often associated with anomalies, and here we show that SIMPS are well-suited to capturing anomalous symmetries. In the present context, an anomalous symmetry can be described as a symmetry which does not permit a short-range entangled symmetric state [Chen2011a]. Given a group of non-onsite symmetries, one can calculate an invariant called a 3-cocycle [Chen2011a, Else2014]. If this 3-cocycle belongs to a non-trivial equivalence class, then it can be shown that no injective MPS can be symmetric under this group [Chen2011a]. The alternatives are (a) the MPS is non-injective, meaning it has long-range order i.e. symmetry breaking, or (b) the state is not representable as an MPS with finite bond dimension, such as critical states of gapless Hamiltonians.
@@ 3.4.1 String order parameters
where $\ell$ and $r$ specify the locations of the left and right ends of the string and $O^{\bf{a}}_{\ell/r}$ ($O^{\bf{b}}_{\ell/r}$) are the aforementioned operators that act in a finite region around sites $\ell/r$ and have the effect of inserting $Z$ ($X$) in the virtual space of the SIMPS. By construction, we have $\langle\psi_{\bf{a},\bf{b}}|S^{{\bf{a}/\bf{b}}}(\ell,r)|\psi_{\bf{a},\bf{b}}\rangle=1$, i.e. perfect string order.
@@ 3.4.2 Entanglement spectrum
There is, however, a hidden entanglement spectrum degeneracy which can be revealed using the SIMPS representation. Consider a SIMPS with OBC (5) and project the spin at site $k$ onto the state $|s\rangle$ (by, e.g. measuring it). This results in the state $|s\rangle_{k}\otimes|\psi_{s}\rangle$ where,
@@ 3.4.4 Universal fingerprints
When then apply Lemma 1 to the gates $u_{i}$. For the gates that act directly on $|\psi_{\bf{a},\bf{b}}\rangle$, i.e. the first layer, we can immediately apply Lemma 1 to replace them with diagonal operators. For the gates in other layers, we need to first commute the gates through the lower layers before we can act on $|\psi_{\bf{a},\bf{b}}\rangle$. This introduces some correlations between the different local operators, but these correlations remain local. By replacing each gate $u_{i}$ with a diagonal operator, we can replace the whole circuit $\mathcal{U}_{\phi}$ with a diagonal operator $\mathcal{J}_{\phi}$ such that $|\phi\rangle=\mathcal{J}_{\phi}|\psi_{\bf{a},\bf{b}}\rangle$. This operator is no longer a finite-depth circuit in general (or even a unitary operator), but it can nonetheless be represented by a matrix product operator (MPO) of the form,
@@ 4.3 Quantum wire in SIMPS
These examples have an important property that is distinct from the MPS examples described in Sec. 3.1. In the cluster state, and more general quantum wires emerging from SPT phases with onsite symmetries, the classical side processing to determine the byproduct operators is a linear function of the measurement outcomes. Conversely, Eq. (53) requires calculating a non-linear boolean function (e.g. $s_{1}s_{2}$). This is significant, as all known schemes of teleportation (and, more generally, MBQC) have linear classical side processing relations [^2] Possible exceptions are given by MBQC schemes that involve an initial round of measurements that reduce the resource state to a cluster state, followed by normal cluster state MBQC [Wei2012, Miller2016]. However, since these protocols are non-deterministic, we do not consider them to be counterexamples.. In fact, this linearity plays a key role in our fundamental understanding of MBQC, such as its relationship to quantum contextuality [Raussendorf2013, Frembs2023]. Therefore, our examples may provide new fundamental insights into more general schemes of MBQC.
@@ 5.1 Variational SIMPS for Rydberg chain
We first consider how such a constrained spin chain can be captured in MPS. Referring to the ground and Rydberg states as $|0\rangle$ and $|1\rangle$, the constraint can implemented by restricting to MPS tensors $A^{i}$ satisfying
@@ 5.2 SIMPS for an exact scar state
Observe that $B^{{\uparrow}{\uparrow}}=B^{{\downarrow}{\downarrow}}=0$ as required by the constraint. Using this form, we can easily deform the AKLT state while preserving the constraint by fixing $B^{{\uparrow}{\uparrow}}=B^{{\downarrow}{\downarrow}}=0$ and freely varying the other components.
@@ 6 Discussion
It was shown in Ref. [Wahl2012] that, by measuring the bulk spins, one can establish a maximally entangled state between a pair of two-dimensional spins that are appropriately coupled to the boundaries. This example initially appears puzzling for a number of reasons. First, there is LRLE despite the fact that the MPS tensors are not unitary, and the amount of entanglement generated is that of a pair of two-dimensional spins rather than three-dimensional as one may have expected from a $D=3$ MPS. Second, it is not clear if the state is a quantum wire, i.e. if there is an efficient classical side processing to determine the byproduct operator. Finally, Ref. [Wahl2012] showed that this state only has a single onsite $Z_{2}$ symmetry, which is not enough to support non-trivial SPT order. Therefore, it is not clear if the LRLE can be understood in terms of SPT order and string order parameters.
@@ 6.3 Future directions
As discussed in Sec. 5, we expect that SIMPS will be useful in numerical studies of constrained spin chains due to the linear implementation of the constraint. However, making use of this constraint would require modifying standard algorithms such that the SIMPS structure is properly taken advantage of (i.e. simply rewriting the SIMPS as an MPS and doing normal MPS algorithms would get rid of the possible advantages). Whether the cost of this modification outweighs the benefits of the linear implementation of the constraint remains an open question. We also hope that SIMPS will be useful for analytical studies of constrained spin chains. For example, they may help to classify phases of matter in constrained spin chains.

### v3

@@ Abstract
We describe a class of spin chains with new physical and computational properties. On the physical side, the spin chains give examples of symmetry-protected topological phases that are defined by non-onsite symmetries, i.e., symmetries that are not a tensor product of single-site operators. These phases can be detected by string-order parameters, but notably do not exhibit entanglement spectrum degeneracy. On the computational side, the spin chains represent a new class of states that can be used to deterministically teleport information across long distances, with the novel property that the necessary classical side processing is a non-linear function of the measurement outcomes. We also give examples of states that can serve as universal resources for measurement-based quantum computation, providing the first examples of such resources without entanglement spectrum degeneracy. The key tool in our analysis is a new kind of tensor network representation which we call split-index matrix product states (SIMPS). We develop the basic formalism of SIMPS, compare them to matrix product states, show how they are better equipped to describe certain kinds of non-onsite symmetries including anomalous symmetries, and discuss how they are also well-suited to describing quantum teleportation and constrained spin chains.
@@ 1 Introduction
In this paper, we construct a class of quantum spin chains that have novel physical and computational properties. The key concept that we introduce is a new kind of tensor network representation we call split-index matrix product states (SIMPS) (4). While equivalent to MPS in terms of which states can be efficiently captured, SIMPS offer an alternative description that we argue is essential to understanding the properties of the constructed models. For example, a key physical property of the models is that they possess symmetries which are non-onsite, i.e., not a tensor product of single-site unitaries. Compared to MPS, we show that SIMPS provide a much simpler way to identify and analyze such symmetries. Similarly, SIMPS offer a clearer picture of the computational capabilities of certain models, including ones that have previously appeared in the literature. We also show that SIMPS are well-suited to capturing constrained spin chains, including models exhibiting quantum many-body scars and Hilbert space fragmentation [Moudgalya2022].
@@ 2 The SIMPS representation
The above definition is equivalent to the condition that the state obtained by projecting every $(L_{1}+1)$-th spin onto an arbitrary fixed basis state is an injective MPS. As in the case of MPS, this property is stable in that, if it holds for some length $L$, then it holds for $L+1$ as well, and therefore for all lengths $>L$. To see this, observe that,
$\mathrm{span}\{B^{s_{1}i_{1}}B^{i_{1}i_{2}}\cdots B^{i_{L+1}s_{2}}\}_{i_{1},\dots,i_{L+1}}$ (6)
$=\mathrm{span}\{B^{s_{1}i_{1}}B^{i_{1}i_{2}}\cdots B^{i_{L}s_{3}}B^{s_{3}s_{2}}\}_{i_{1},\dots,i_{L};s_{3}}$
$\supset\mathrm{span}\{B^{s_{1}i_{1}}B^{i_{1}i_{2}}\cdots B^{i_{L-1}s_{3}}B^{s_{3}s_{2}}\}_{i_{1},\dots,i_{L-1};s_{3}}$
$=\mathrm{span}\{B^{s_{1}i_{1}}B^{i_{1}i_{2}}\cdots B^{i_{L}s_{2}}\}_{i_{1},\dots,i_{L}}$
where, going from the second line to the third line, we used the normality for length $L$, which holds for all fixed values of $s_{3}$.
As with MPS, we can construct an operator that reveals the virtual space of a normal SIMPS. Define the map,
@@
For $L\geq L_{1}$, this map is invertible and its inverse $B_{L}^{-1}$ acts on $L+2$ spins and serves the purpose of revealing the virtual indices of the SIMPS. Later, we will use this map to construct string order parameters. One could also use this map to construct parent Hamiltonians for SIMPS which may have different properties than they corresponding MPS parent Hamiltonians, but we do not pursue this direction here.
@@ 2.1 Converting between MPS and SIMPS
First, suppose we have a SIMPS tensor $B^{ij}$. We can convert the SIMPS to an MPS by cutting $B^{ij}$ in half via singular value decomposition (SVD). Namely, define $\mathbb{B}=\sum_{i,j,\alpha,\beta}B^{ij}_{\alpha\beta}|i\alpha\rangle\langle j\beta|$ which is a $\chi\times\chi$ matrix with $\chi=\sum_{i}\chi_{i}$ and use the SVD to write $\mathbb{B}=usv^{\dagger}$ where $u,v$ are isometries mapping from $\mathbb{C}^{\chi}$ to $\mathbb{C}^{D}$ satisfying $u^{\dagger}u=v^{\dagger}v=\mathbb{1}$ and $s$ is a $D\times D$ diagonal matrix. Therein, $D$ is the rank of $\mathbb{B}$. For compactness, let us absorb $s$ into $v$, defining $w^{\dagger}=sv^{\dagger}$. From this, we define the matrices $A^{i}$ as,
@@
In the generic case, $A^{i}$ is full rank so $P^{i}=\mathbb{1}$ and $B^{ij}=A^{i}$. The cases where the SIMPS representation is distinct from the MPS representation are therefore those for which $A^{i}$ is not full-rank in some physical basis $\{i\}$. This leads us to the following definition of when a MPS “makes a good SIMPS”; i.e., when the SIMPS and MPS representations differ,
@@ 2.2 Fundamental theorem of SIMPS
We now state the SIMPS equivalent of the fundamental theorem of MPS. The theorem describes how two normal SIMPS tensors generating the same state are related. For MPS, the fundamental theorem states that if $A^{i}$ and $B^{i}$ are both normal and generate the same MPS, then there is an invertible matrix $V$ such that $A^{i}=VB^{i}V^{-1}$, i.e., the two tensors are related by a so-called gauge transformation [Perez-Garcia2008]. For injective SIMPS, gauge transformations like $B^{ij}\rightarrow VB^{ij}V^{-1}$ can still be used to obtain new tensors generating the same state. However, it turns out that a larger class of transformations is also possible,
@@ 2.3 Examples of SIMPS
Example 1: The cluster state. First, we consider the 1D cluster state, which is a paradigmatic resource state for MBQC [Raussendorf2001a], and also a simple example of a state with SPT order [Son2012]. The cluster state can be written as,
@@
It is straightforward to verify that this MPS is normal, so it describes the unique ground state of a gapped Hamiltonian. Both matrices are rank deficient, having rank equal to two. Therefore, using Eq. (10), we can derive the corresponding SIMPS tensor with bond dimension $\chi=2$,
@@ 3 Non-onsite symmetries and SPT phases in SIMPS
In this section, we demonstrate the one context where SIMPS are advantageous over MPS: capturing non-onsite symmetries. First, we review how onsite symmetries are captured in MPS. Then, we show how similar ideas can be used to analyze certain non-onsite symmetries in SIMPS. This includes a general class of anomalous group-like symmetries. We then define a class of SIMPS possessing non-trivial SPT order protected by non-onsite symmetries and characterize their physical properties.
@@ 3.1 SPT order in MPS
If there is a set of symmetries $u_{g}$ that form a representation of a group $G$ such that $u_{g}u_{h}=u_{gh}$, then the corresponding operators $V_{g}$ that appear in Eq. (22) form a projective representation, $V_{g}V_{h}=\omega(g,h)V_{gh}$ for some phase factors $\omega(g,h)$ [Schuch2011]. If $G$ is finite abelian, which we always assume from now on, we can instead characterize the representation by the relation $V_{g}V_{h}=\Omega(g,h)V_{h}V_{g}$ where $\Omega(g,h)=\omega(g,h)\omega(h,g)^{-1}$. In the case of finite abelian $G$, the phases $\Omega(g,h)$ uniquely label the SPT phase to which the MPS belongs [Berkovich1998]. If these phases are not all equal to one, then the MPS belongs to a non-trivial SPT phase protected by $G$ [Pollmann2010, Chen2011, Schuch2011].
@@ 3.3 Anomalous symmetries
Non-onsite symmetries are often associated with anomalies, and here we show that SIMPS are well-suited to capturing anomalous symmetries. In the present context, an anomalous symmetry can be described as a symmetry which does not permit a short-range entangled symmetric state [Chen2011a]. Given a group of non-onsite symmetries, one can calculate an invariant called a 3-cocycle [Chen2011a, Else2014]. If this 3-cocycle belongs to a non-trivial equivalence class, then it can be shown that no injective MPS can be symmetric under this group [Chen2011a]. The alternatives are (a) the MPS is non-injective, meaning it has long-range order i.e., symmetry breaking, or (b) the state is not representable as an MPS with finite bond dimension, such as critical states of gapless Hamiltonians.
@@ 3.4.1 String order parameters
where $\ell$ and $r$ specify the locations of the left and right ends of the string and $O^{\bf{a}}_{\ell/r}$ ($O^{\bf{b}}_{\ell/r}$) are the aforementioned operators that act in a finite region around sites $\ell/r$ and have the effect of inserting $Z$ ($X$) in the virtual space of the SIMPS. By construction, we have $\langle\psi_{\bf{a},\bf{b}}|S^{{\bf{a}/\bf{b}}}(\ell,r)|\psi_{\bf{a},\bf{b}}\rangle=1$, i.e., perfect string order.
@@ 3.4.2 Entanglement spectrum
There is, however, a hidden entanglement spectrum degeneracy which can be revealed using the SIMPS representation. Consider a SIMPS with OBC (5) and project the spin at site $k$ onto the state $|s\rangle$ (by, e.g., measuring it). This results in the state $|s\rangle_{k}\otimes|\psi_{s}\rangle$ where,
@@ 3.4.4 Universal fingerprints
When then apply Lemma 1 to the gates $u_{i}$. For the gates that act directly on $|\psi_{\bf{a},\bf{b}}\rangle$, i.e., the first layer, we can immediately apply Lemma 1 to replace them with diagonal operators. For the gates in other layers, we need to first commute the gates through the lower layers before we can act on $|\psi_{\bf{a},\bf{b}}\rangle$. This introduces some correlations between the different local operators, but these correlations remain local. By replacing each gate $u_{i}$ with a diagonal operator, we can replace the whole circuit $\mathcal{U}_{\phi}$ with a diagonal operator $\mathcal{J}_{\phi}$ such that $|\phi\rangle=\mathcal{J}_{\phi}|\psi_{\bf{a},\bf{b}}\rangle$. This operator is no longer a finite-depth circuit in general (or even a unitary operator), but it can nonetheless be represented by a matrix product operator (MPO) of the form,
@@ 4.3 Quantum wire in SIMPS
These examples have an important property that is distinct from the MPS examples described in Sec. 3.1. In the cluster state, and more general quantum wires emerging from SPT phases with onsite symmetries, the classical side processing to determine the byproduct operators is a linear function of the measurement outcomes. Conversely, Eq. (54) requires calculating a non-linear boolean function (e.g., $s_{1}s_{2}$). This is significant, as all known schemes of teleportation (and, more generally, MBQC) have linear classical side processing relations [^2] Possible exceptions are given by MBQC schemes that involve an initial round of measurements that reduce the resource state to a cluster state, followed by normal cluster state MBQC [Wei2012, Miller2016]. However, since these protocols are non-deterministic, we do not consider them to be counterexamples.. In fact, this linearity plays a key role in our fundamental understanding of MBQC, such as its relationship to quantum contextuality [Raussendorf2013, Frembs2023]. Therefore, our examples may provide new fundamental insights into more general schemes of MBQC.
@@ 5.1 Variational SIMPS for Rydberg chain
We first consider how such a constrained spin chain can be captured in MPS. Referring to the ground and Rydberg states as $|0\rangle$ and $|1\rangle$, the constraint can be implemented by restricting to MPS tensors $A^{i}$ satisfying
@@ 5.2 SIMPS for an exact scar state
Observe that $B^{{\uparrow}{\uparrow}}=B^{{\downarrow}{\downarrow}}=0$ as required by the constraint. Using this form, we can easily deform the AKLT state while preserving the constraint by fixing $B^{{\uparrow}{\uparrow}}=B^{{\downarrow}{\downarrow}}=0$ and freely varying the other components. This could facilitate, e.g., finding an exact path of SIMPS in the constrained Hilbert space that goes through a phase transition, similar to the MPS paths derived in Refs. [Wolf2006, Jones2021].
@@ 6 Discussion
It was shown in Ref. [Wahl2012] that, by measuring the bulk spins, one can establish a maximally entangled state between a pair of two-dimensional spins that are appropriately coupled to the boundaries. This example initially appears puzzling for a number of reasons. First, there is LRLE despite the fact that the MPS tensors are not unitary, and the amount of entanglement generated is that of a pair of two-dimensional spins rather than three-dimensional as one may have expected from a $D=3$ MPS. Second, it is not clear if the state is a quantum wire, i.e., if there is an efficient classical side processing to determine the byproduct operator. Finally, Ref. [Wahl2012] showed that this state only has a single onsite $Z_{2}$ symmetry, which is not enough to support non-trivial SPT order. Therefore, it is not clear if the LRLE can be understood in terms of SPT order and string order parameters.
@@ 6.3 Future directions
As discussed in Sec. 5, we expect that SIMPS will be useful in numerical studies of constrained spin chains due to the linear implementation of the constraint. However, making use of this constraint would require modifying standard algorithms such that the SIMPS structure is properly taken advantage of (i.e., simply rewriting the SIMPS as an MPS and doing normal MPS algorithms would get rid of the possible advantages). Whether the cost of this modification outweighs the benefits of the linear implementation of the constraint remains an open question. We also hope that SIMPS will be useful for analytical studies of constrained spin chains. For example, they may help to classify phases of matter in constrained spin chains.